The mean of the ages of father and his son is 27 years. After 18 years, father will be twice as old as his son. Their present ages are
Answer by Horlaye(27)

Let the father and son's age be represented by x and y respectively
From the 1st sentence
(x+y)/2 = 27
Cross multiply the equation
x+y = 2*27
x+y = 54..........(1)
From the 2nd sentence
x+18 = 2(y+18)
x+18 = 2y+36
x-2y = 36-18
x-2y = 18..........(2)
Bring the two equations together
x+y = 54..........(3)
x-2y = 18..........(4)
Eliminate equation (4) from equation (3)
3y = 36
y = 36/3
y = 12
Substitute y = 12 in equation (3)
x+y = 54
x+12 = 54
x = 54-12
x = 42
Therefore there present ages =
Father = x = 42
Son = y = 12